OpenStudy (anonymous):

You shade four squares in a grid. How many squares are there is the shaded portion represents 20 percent of the grid?

OpenStudy (blurbendy):

20. 4/20 = 1/5 1/5 = 20 percent.

OpenStudy (anonymous):

Ok thanks! :D So 20 squares on the grid?

OpenStudy (blurbendy):

Yup! 20 squares total.